# The SAGES Critical View of Safety Challenge: A Global Benchmark for AI-Assisted Surgical Quality Assessment

### v1

@@ Abstract
journal: arXiv Preprint
@@ 4.3 Challenge Tasks and Evaluation Metrics
Participants were evaluated on a held-out test set of 300 laparoscopic cholecystectomy videos submitted through the Grand Challenge platform using Docker-based containers. Test videos and annotations were withheld, and all models were required to operate causally using only past and current frames at inference time. Each submission was evaluated across three subchallenges designed to probe different facets of the modeling problem.
@@ 4.5 Awards
Monetary prizes were awarded to the top 3 ranking teams for each of the 3 subchallenges. The overall winner across the 3 subchallenges was additionally awarded an NVIDIA IGX edge AI platform.
@@ 5.1 Baseline: LG-DG
5 SRV-WEISS University College London (UK) EndoViT + Dense Prediction Transformer, EfficientNet, ConvNeXt Segmentation + classification fusion None Self-Supervised (EndoViT) + Supervised (EndoViT on Endoscapes segmentation; EfficientNet on Endoscapes classes)
6 Ostrich Anonymous DenseNet-121 (5× ensemble) None None Supervised (ImageNet weights)
@@ 5.2.5 Team SRV-WEISS: Hybrid Segmentation–Classification Pipeline for CVS Prediction
To adapt to the expanded input dimensionality caused by concatenated segmentations, the ConvNeXt input layer was modified using a weight averaging strategy [66]. All predictions were made per-frame, and training was fully supervised using a weighted binary cross-entropy loss. The weighting was conditioned on the prevalence of each class in the training data. The DPT weights were frozen during training to reduce overfitting and stabilize learning.
5.2.6 Team Ostrich: DenseNet Ensemble for Multi-Criteria CVS Prediction
@@ 5.3 Comparative analysis of methods
    Standard general pretraining only: Rely on generic computer vision dataset (e.g., ImageNet) pretraining, without additional surgical datasets. Teams: theator, IRCV-URV, HFUT-MedIA, Ostrich, CVS HUST, Pandas.
@@
      Detection pretraining on Endoscapes-BBox201: mmll (Mask R-CNN detector).
@@
    CNN: FightTumor (ConvNeXt-B), Pandas (ConvNeXt V2), IRCV-URV (EfficientNet-B5 + Feature Pyramid Network with spatial attention), Ostrich (DenseNet-121).
@@
    Used ensembling: mmll (3 heterogeneous graph- and transformer-based models), Ostrich (5 DenseNet-121), Theator (6 EVA-02-ViT-L via model soups), Farm (7 varied backbones), TUE-VCA (5× PVT-v2).
@@
    Classification only: Ostrich, FightTumor, HFUT-Media, IRCV-URV, Farm
@@ 6.1 Overview
FightTumor 40.85 76.34 47.15 54.78 TUE-VCA 0.059 0.050 0.046 0.052 Ostrich 19.72 63.79 16.82 49.24
Ostrich 40.35 75.32 44.13 53.27 Farm 0.063 0.057 0.055 0.058 FightTumor 27.81 63.22 30.91 42.83
SRV-WEISS 38.16 66.27 42.29 48.91 Ostrich 0.092 0.091 0.077 0.086 IRCV-URV 26.00 59.83 14.32 41.38
@@ 7 Limitations
This manuscript concentrates on methodological trends across submissions rather than a full data-centric analysis of performance across all dataset attributes. The dataset and manual quality assessments enable several high-value directions. First, the quality assessments themselves constitute one of the largest video-based initiatives in surgery, even without the AI component. Mining these assessments for predictors of difficult frames, and using those signals to trigger targeted model development, could directly benefit clinical reliability. Second, the data support stratified analyses that each warrant independent investigation, for example performance in procedures from LMIC settings, under-represented countries, or other cohort characteristics. In this work we prioritised a single coherent synthesis of methodological trends to provide clear takeaways the community can build on.
@@ Acknowledgment
We would like to thank Surgical Safety Technologies for their support in enabling the video collection and de-identification infrastructure. We would like to thank all contributing members to each of the submissions, including Leon Mayer, Patrick Godau, Piotr Kalinowski, Lars Krämer, Ömer Sümer, Dominik Michael, Tim Rädsch, Fabian Isensee, Lena Maier-Hein, Sanket Rajan Gupte, Elaine Sui, Alan Brown, Josiah Aklilu, Reid Dale, Charlotte Egeland, Jeffrey Heo, Matthew Leipzip, Anita Rau, Joshua Villarreal, Jeffrey Jopling, Dan Azagury, Serena Yeung-Levy, Pranav Poudel, Sachin Acharya, Sophia Bano, Euan Sexton, Yuanbin Wang, Doménec Savi Puig Valls, Chengzhi Hu, Zhiwei Wang, Yuxuan Yang, Rui Xu, Shuai Ding.
@@
This work was supported by French state funds managed by the ANR under Grant ANR-20-CHIA-0029-01 (Chair AI4ORSafety) and Grant ANR-10-IAHU-02 (IHU Strasbourg), and by BPI France under reference DOS0180017/00 (project 5G-OR). This work was partially supported by a CRICO grant. Awards for the challenge were sponsored by NVIDIA, Olympus, Intuitive and Medtronic plc.
Participating teams would like to acknowledge the following funding: Farm: Wellcome Leap SDS-HD: National Center for Tumor Diseases (NCT) Heidelberg, Helmholtz Imaging SRV-WEISS: EPSRC
CRediT authorship contribution statement
DA : Conceptualization, Data Curation, Data Analysis and Interpretation, Methodology, Software, Investigation, Validation, Evaluation, Formal Analysis, Visualization, Writing - Original Draft, Writing - Review & Editing, Challenge Organization, Resources. DN, OB, AY, JH, XW, RL, LL, YW, SK, BB, TJ, ZM, AW, JM, YX, ZW, AMOA, HAFR, BS, KY, YZ, HW : Methodology, Software, Writing - Review & Editing. PI : Conceptualization, Writing - Review & Editing, Supervision, Challenge Organization, Resources, Funding Acquisition, Project Administration.

### v2

@@ Abstract
journal: arXiv Preprintfn3fn3footnotetext: The SAGES CVS Challenge Data Consortium is represented by Brian Quaranto, Mika Makari, Omero Costa Filho, Giovanni Taffurelli, Alexander Farrell, Taner Shakir, Henry Badgery, Alexandros Chamzin, Fernando Ponce Leon, Salvador Morales-Conde, Lovenish Bains, Benjamin Babic, Micheli Domingos, Yuri Fishman, Ankit Patel.
@@ 4.3 Challenge Tasks and Evaluation Metrics
Participants were evaluated on a held-out test set of 300 laparoscopic cholecystectomy videos submitted through the Grand Challenge platform using Docker-based containers. Test videos and annotations were withheld, and all models were required to operate causally using only past and current frames at inference time. Each submission was evaluated across three subchallenges designed to probe different facets of the modeling problem. General updates regarding the challenge were made available through the challenge website  [^6] https://www.cvschallenge.org/. Members of the organizing team and their immediate research groups were not permitted to participate in the Challenge.
@@ 4.5 Awards
Monetary prizes were awarded to the top 3 ranking teams for each of the 3 subchallenges and announced at MICCAI 2024. The overall winner across the 3 subchallenges was additionally awarded an NVIDIA IGX edge AI platform.
@@ 5.1 Baseline: LG-DG
5 SRV-WEISS University College London (UK) EndoViT + Dense Prediction Transformer, EfficientNet, ConvNeXt None Segmentation + classification fusion Self-Supervised (EndoViT) + Supervised (EndoViT on Endoscapes segmentation; EfficientNet on Endoscapes classes)
6 Caresyntax Caresyntax DenseNet-121 (5× ensemble) None None Supervised (ImageNet weights)
@@ 5.2.5 Team SRV-WEISS: Hybrid Segmentation–Classification Pipeline for CVS Prediction
To adapt to the expanded input dimensionality caused by concatenated segmentations, the ConvNeXt input layer was modified using a weight averaging strategy [66]. All predictions were made per-frame, and training was fully supervised using a weighted binary cross-entropy loss. The weighting accounted for both the prevalence of each class and the annotator’s confidence, i.e., an annotator-confidence-aware label distribution. The DPT weights were frozen during training to reduce overfitting and stabilize learning.
5.2.6 Team Caresyntax: DenseNet Ensemble for Multi-Criteria CVS Prediction
@@ 5.3 Comparative analysis of methods
    Standard general pretraining only: Rely on generic computer vision dataset (e.g., ImageNet) pretraining, without additional surgical datasets. Teams: theator, IRCV-URV, HFUT-MedIA, Caresyntax, CVS HUST, Pandas.
@@
      Detection pretraining on Endoscapes-BBox201: mmll (Mask R-CNN detector), SRV-WEISS (EfficientNet).
@@
    CNN: FightTumor (ConvNeXt-B), Pandas (ConvNeXt V2), IRCV-URV (EfficientNet-B5 + Feature Pyramid Network with spatial attention), Caresyntax (DenseNet-121).
@@
    Used ensembling: mmll (3 heterogeneous graph- and transformer-based models), Caresyntax (5 DenseNet-121), Theator (6 EVA-02-ViT-L via model soups), Farm (7 varied backbones), TUE-VCA (5× PVT-v2).
@@
    Classification only: Caresyntax, FightTumor, HFUT-Media, IRCV-URV, Farm
@@ 6.1 Overview
FightTumor 40.85 76.34 47.15 54.78 TUE-VCA 0.059 0.050 0.046 0.052 Caresyntax 19.72 63.79 16.82 49.24
Caresyntax 40.35 75.32 44.13 53.27 Farm 0.063 0.057 0.055 0.058 FightTumor 27.81 63.22 30.91 42.83
SRV-WEISS 38.16 66.27 42.29 48.91 Caresyntax 0.092 0.091 0.077 0.086 IRCV-URV 26.00 59.83 14.32 41.38
@@ 7 Limitations
This manuscript, following the BIAS guidelines [maier2020bias], concentrates on methodological trends across submissions rather than a full data-centric analysis of performance across all dataset attributes. The dataset and manual quality assessments enable several high-value directions. First, the quality assessments themselves constitute one of the largest video-based initiatives in surgery, even without the AI component. Mining these assessments for predictors of difficult frames, and using those signals to trigger targeted model development, could directly benefit clinical reliability. Second, the data support stratified analyses that each warrant independent investigation, for example performance in procedures from LMIC settings, under-represented countries, or other cohort characteristics. In this work we prioritised a single coherent synthesis of methodological trends to provide clear takeaways the community can build on.
@@ Acknowledgment
We would like to thank Surgical Safety Technologies for their support in enabling the video collection and de-identification infrastructure. We would like to thank all contributing members to each of the submissions, including Leon Mayer, Patrick Godau, Piotr Kalinowski, Lars Krämer, Ömer Sümer, Dominik Michael, Tim Rädsch, Fabian Isensee, Lena Maier-Hein, Sanket Rajan Gupte, Elaine Sui, Alan Brown, Josiah Aklilu, Reid Dale, Charlotte Egeland, Jeffrey Heo, Matthew Leipzip, Anita Rau, Joshua Villarreal, Jeffrey Jopling, Dan Azagury, Serena Yeung-Levy, Pranav Poudel, Sachin Acharya, Sophia Bano, Euan Sexton, Yuanbin Wang, Doménec Puig, Chengzhi Hu, Zhiwei Wang, Yuxuan Yang, Rui Xu, Shuai Ding.
@@
This work was supported by French state funds managed by the ANR under Grant ANR-20-CHIA-0029-01 (Chair AI4ORSafety) and Grant ANR-10-IAHU-02 (IHU Strasbourg), and by BPI France under reference DOS0180017/00 (project 5G-OR). This work was also developed within the Interdisciplinary Thematic Institute HealthTech (ITI 2021-2028 program of the University of Strasbourg, CNRS and Inserm), supported by IdEx Unistra (ANR-10-IDEX-0002) and SFRI (STRATUS project, ANR-20-SFRI-0012) under the framework of the French Investments for the Future Program. This work has received funding from the European Union (ERC, CompSURG, 101088553). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was also partially supported by a CRICO grant. Awards for the challenge were sponsored by NVIDIA, Olympus, Intuitive and Medtronic plc.
Participating teams would like to acknowledge the following funding: Farm: Wellcome Leap SDS-HD: National Center for Tumor Diseases (NCT) Heidelberg, Helmholtz Imaging SRV-WEISS: EPSRC (EP/Y01958X/1,EP/W00805X/1)
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT (OpenAI) to refine language, improve readability, and ensure clarity of expression. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the final version of the manuscript.
Data and code availability
The dataset used in this study will be made available. Benchmark evaluation code for metric calculation is publicly accessible within the challenge repository [^7] https://github.com/SAIIL/CVS_challenge_code/tree/main.